OpenStudy (loser66):

Suppose that the length of time Y takes a worker to complete a certain task has the probability density function given by \(f(y) = \begin{cases} e^{-(y-\theta)}~~~if y>\theta \\0~~~elsewhere\end{cases}\) where \(\theta\) is a positive constant that represents the minimum time until task completion. Let \(Y_1,Y_2,...,Y_n\) denote a random sample of completion times from this distribution. Find a)The density function for \(Y_(1)=min(Y_1,Y_2,...,Y_n)\) b) \(E(Y_(1))\) Please, help

OpenStudy (loser66):

@Zarkon

OpenStudy (loser66):

\[f(y)=\begin{cases} e^{-(y-\theta)}~~~if ~~~y>\theta\\0~~~elsewhere \end{cases}\]

OpenStudy (loser66):

\(f_Y(Y_{(1)})=n (1-F_Y(y)^{n-1})f_Y(y)\)

OpenStudy (zarkon):

closer

OpenStudy (loser66):

I have just that formula.

OpenStudy (loser66):

I don't know how to get closer. :)

OpenStudy (zarkon):

\[f_{Y_{(1)}}(y)=n [1-F_Y(y)]^{n-1}f_Y(y)\]

OpenStudy (loser66):

So, what is difference between \(f_{Y_1}(y)\) and \(f_Y(y_{(1)})\)

OpenStudy (zarkon):

one is correct notation and one is not

OpenStudy (zarkon):

also notice how the n-1 power go to the entire quantity

OpenStudy (loser66):

ok. :) My question is how to find \(F_Y(y)\)? if I take integrating of \(f_Y(y)\), what are the limits?

OpenStudy (zarkon):

\(\theta\) to \(y\)

OpenStudy (loser66):

why? why is it not from theta to infinitive?

OpenStudy (zarkon):

if you did that you would get 1

OpenStudy (zarkon):

\[F_Y(y)=P(Y\le y)=\int_{-\infty}^{y}f_Y(t)dt\]

OpenStudy (loser66):

I got it.

OpenStudy (loser66):

if I do b), I have to apply formula \(E(Y(1))=\int(y f_{Y_1} (y)dy\) right? what are the limits? same as part a?

OpenStudy (zarkon):

there you integrate over the entire support

OpenStudy (loser66):

I don't get what you mean.

OpenStudy (zarkon):

from theta to infinity

OpenStudy (loser66):

Is there anyway else or just that way?

OpenStudy (zarkon):

i can't think of another way of the top of my head

OpenStudy (loser66):

We have N (normal), S (sampling), T (just call t), chi-square what is Z,?

OpenStudy (zarkon):

Z the random variable?

OpenStudy (loser66):

Yes

OpenStudy (zarkon):

typically Z is a standard normal random variable

OpenStudy (loser66):

you mean Z~N(0,1)?

OpenStudy (zarkon):

yes

OpenStudy (loser66):

On the formula, n, n-1 apply to find some value of distribution, but they are number of freedom, right? if n large enough, how are they different to each other?

OpenStudy (zarkon):

n just represents the number of random variable you are considering.

OpenStudy (zarkon):

at least for this problem

OpenStudy (loser66):

I am ok with this problem. Thanks for the help. I just want to make clear the notations from my lecture. :)

OpenStudy (zarkon):

did you find the expected value?

OpenStudy (loser66):

Sometimes, my Prof used U, V, Z, W, T, S..... some of them are universal notation, some are just the name of a particular problem. I got confused.

OpenStudy (zarkon):

ah...ic

OpenStudy (loser66):

I didn't find it yet.

OpenStudy (zarkon):

ok...I did...if you get an answer let me know

OpenStudy (loser66):

I know you are a statistic Prof so that I try to ask my unclear information first. :)

OpenStudy (loser66):

Yes, I will. Again, thanks for the help

OpenStudy (loser66):

I am sorry. I was on the phone. This is my work

OpenStudy (zarkon):

good

OpenStudy (loser66):

:)